### v1

# Pre-training Large Memory Language Models with Internal and External Knowledge

Abstract
Neural language models are black-boxes – both linguistic patterns and factual knowledge are distributed across billions of opaque parameters. This entangled encoding makes it difficult to reliably inspect, verify, or update specific facts. We propose a new class of language models, Large Memory Language Models (LMLM) [^2] LmLm stands for Large Memory Language Model and is pronounced “LamLam”. with a pre-training recipe that stores factual knowledge in both internal weights and an external database. Our approach strategically masks externally retrieved factual values from the training loss, thereby teaching the model to perform targeted lookups rather than relying on memorization in model weights. Our experiments demonstrate that LmLms achieve competitive performance compared to significantly larger, knowledge-dense LLMs on standard benchmarks, while offering the advantages of explicit, editable, and verifiable knowledge bases. This work represents a fundamental shift in how language models interact with and manage factual knowledge. 33footnotetext: We will open-source our code and models at https://github.com/kilian-group/LMLM.
1 Introduction
Large language models (LLMs) encode very different capabilities in their neural weights: vast factual knowledge and linguistic competency. These capabilities are intertwined inside the model, with significant implications for both training and inference. For instance, facts need to be observed many times during training to be reliably remembered, and it is difficult to forget specific facts after training. This tight coupling of knowledge and linguistic ability makes updating one without affecting the other extremely challenging. As an alternative, related work have studied offloading knowledge, such as nearest-neighbor language models that maintain access to the entire training corpus [27, 30]. However, these works do not distinguish between different types of knowledge and require constant, indiscriminate look-ups.
We take a middle ground between storing knowledge internally or externally: instead of retrieving for every token, we explicitly retrieve facts, and allow the language model to learn when to best leverage its internal linguistic knowledge versus when to access an external knowledge base.
Consider the sentence “Napoleon was born on August 15th, 1769 in Ajaccio, Corsica to Letizia Bonaparte” (Figure 1). The common knowledge here includes general facts: that Napoleon had a mother who gave birth to him at a certain place and time. In contrast, the specific knowledge includes the exact date, location, and names. This distinction is not merely about frequency. Common knowledge generalizes across contexts, making it well-suited for storage in the model’s parameters. Specific knowledge is non-generalizable: it appears in narrow contexts, does not transfer, and contributes disproportionately to the memorization burden. Recognizing this asymmetry allows us to design models that are more parameter-efficient, verifiable, and editable.
Figure 1: Spectrum of knowledge representation in pre-training, from fully parametric (e.g., LLMs) to fully externalized (e.g., retrieval from a datastore). Our Large Memory Language Models (LmLms) are designed to both internalize knowledge and retrieve it from external sources.
We introduce Large Memory Language Models (LmLm), a new class of language models with an externalized database aiming to decouple the memorization of specific knowledge from the model weights. We provide an integrated solution for LmLms spanning data preparation, pre-training, and inference. To prepare the training data, we annotate the pre-training corpus with database lookups to offload factual content, using a small, cost-effective LM, fine-tuned for this task. During pre-training, the returned facts are masked from the loss, systematically separating factual knowledge from the neural weights. During inference, instead of recalling memorized facts, the model queries the database.
We compare our LmLms against their LLM counterparts of the same size, pre-trained on the same amount of data without external memory. We find that smaller LmLms achieve comparable performance to LLMs much larger in size on both standard language modeling tasks as well as factual precision on FactScore [36] and T-REx [45]. For instance, LmLm with 382M parameters achieves comparable performance to LLaMA2-7B on these factual precision benchmarks. Additionally, a natural benefit of decoupling knowledge from model parameters is that editing or unlearning becomes a simple operation on the external memory. To verify this, we demonstrate that LmLm can achieve instant forgetting without compromising the general capabilities of the model on TOFU machine unlearning benchmark [33].
While our current experiments are constrained to relatively modest datasets and model scales due to computational limitations, our results reveal compelling trends that suggest LmLms offer a pathway toward language models with substantially reduced dependency on parameter count for factual accuracy. LmLms have the potential to be integrated with techniques from broad fields, including knowledge representation [39], knowledge editing [59], symbolic reasoning [8], as well as mechanistic interpretability [1]. This approach presents promising opportunities for future research to capitalize on the real-time, verifiable knowledge update capabilities inherent to external knowledge databases—potentially transforming how language models store, access, and maintain knowledge.
Figure 2: Overview of the LmLm framework. Our framework consists of  (Left) Data Preparation, where entity-level facts are automatically annotated and stored in an external database;  (Middle) Pretraining, where the model is trained on the annotated text while excluding return values from the loss to discourage memorization; and  (Right) Inference, where the model interleaves text generation with database lookups to ground its outputs on retrieved facts.
2 Large Memory Language Models
LmLm is a new class of language models designed to offload factual knowledge to an external database rather than store it implicitly in model parameters (Figure 2). We first describe the automated annotation during pre-training data preparation (Section 2.1), then discuss pre-training with lookup masking (Section 2.2), and finally how we perform inference with LmLms (Section 2.3).
2.1 Data Preparation: Extracting Knowledge
We begin by extracting atomic entity-level factual knowledge from text and constructing a compact external database. These extracted facts are then interleaved with the original training corpus through explicit lookup calls.
Knowledge Specification. We primarily focus on entity-level atomic factual knowledge, a natural starting point within the broader LmLm framework. We define facts as triplets of the form: (entity, relation) $\to$ value. This level of granularity aligns with previous definitions [34] and represents the most compact and tractable form of factual knowledge to disentangle from the intertwined linguistic patterns in raw text. It also naturally maps to a knowledge graph structure, where triplets define nodes and edges [9]. These atomic facts are ideal for externalization: they are straightforward to extract and verify, yet hard to encode in the model parameters, making them well-suited for storage in an external database.
Automating Knowledge Extraction. Manually extracting factual triples and constructing knowledge graphs at the scale of pre-training data is a major challenge [21]. Additionally, the choice of which facts should be externalized and when the model should issue a database lookup is underexplored. To address this, we distill high-quality annotations from GPT-4o into a lightweight Annotator model that can scalably identify and externalize factual knowledge from raw text, through the following (Figure 3):
  • 1.
    Seed annotation: We use GPT-4o to annotate a small seed dataset of $M$ knowledge-intensive documents with lookup calls and return values. We use $M=1000$ in our setting.
  • 2.
    Filtering: We fine-tune a Corrector model (LLaMA-3.1-8B-Instruct) on the seed annotations to adapt it to the database lookup format. We intentionally underfit the model to identify and filter out lookup calls that are contextually unsupported, overly specific, or improperly formatted. This results in a clean dataset containing non-trivial facts suitable for externalization.
  • 3.
    Annotation: We apply instruction-tuning to an Annotator model (also LLaMA-3.1-8B-Instruct) on the cleaned data and use it to annotate the full pre-training corpus at scale. See Appendix A.1 for additional details.
This annotation process serves two purposes: (1) Database Construction: The extracted triplets form a token-efficient external database that scales with the size of the pre-training corpus. (2) Pre-training Corpus Generation: Lookup calls are interleaved with the original text, enabling the model to learn when to rely on internal knowledge and when to issue a lookup.
Figure 3: Training the Annotator model. We distill high-quality annotations from GPT-4o into a lightweight model that learns to identify and externalize factual knowledge from raw pre-training text, enabling scalable annotation of the full corpus.
2.2 Pre-training
We adopt a standard next-token prediction setup with one critical modification: During pre-training, tokens corresponding to the retrieved factual values are excluded from the loss computation (see Appendix A.3).
This design discourages LmLms from memorizing facts that are offloaded to the external database. Instead, the model learns to ground its generation on retrieved lookup values, reducing the entropy of upcoming tokens, and making training more efficient. This effect is supported by our empirical observation of lower validation perplexity throughout training (see Section 3.2). Intuitively, when the model can rely on accurate externally provided facts, it no longer needs to expand capacity learning complex long-tail distributions of factual content. This insight is consistent with Rho-1 [31], which showed that not all tokens are beneficial to pre-training, and that focusing on informative tokens during training leads to better models.
2.3 Inference
During inference, LmLm generates text autoregressively until a special token is produced, which triggers a database lookup. The model then generates a query of the form “lookup (entityrelation)$\to$”  [^1] Special tokens are used to format the query and retrieved value. and retrieves the corresponding value from the database. The retrieved value is then appended to the context, and the model resumes generation. To perform retrieval, we use fuzzy matching of the query in the all-MiniLM-L6-v2 sentence embedding space. Additional details are provided in Appendix A.4.
Figure 4: Results overview. (Left) Normalized validation perplexity throughout training: LmLm achieves lower perplexity, indicating that offloading factual knowledge improves training efficiency. (Middle) FactScore vs. average natural language understanding (NLU) performance: LmLm significantly improves factual precision over its Standard counterparts without sacrificing NLU performance. (Right) Machine unlearning: LmLm forgets targeted facts while preserving general model performance.
3 Experiments
We evaluate LmLms via validation perplexity (Section  3.2), factual precision (Section  3.3), as well as machine unlearning (Section  3.4), with summary of these results in Figure 4. We further discuss additional implications on LmLm in Section 4.
3.1 Experimental Setup
Pretraining and Model Setup. We pretrain on a high-quality Wikipedia corpus ($\sim\!3\text{B}$ tokens) from the OLMo2 project [^2] https://huggingface.co/datasets/allenai/dolmino-mix-1124 [18], and evaluate perplexity on a held-out set of 1,000 samples ($\sim\!245\text{k}$ tokens). We pre-train LmLm from scratch using GPT-2 and LLaMA2-style architectures with their standard tokenizers and vocabularies, extended by four special tokens for lookup calls. All models are trained for 8 epochs with a context length of 1,024-tokens, using mixed precision. For LLaMA2-176M and LLaMA2-382M, we use a batch size of 256 over 105k steps. Training completes within 8 H100-days. Full configuration details are in Appendix A.2.
Database and Retrieval Setting. We construct the database by annotating the entire pretraining corpus, resulting in 54.6M knowledge triplets. Retrieval uses fuzzy matching with cosine similarity over all-MiniLM-L6-v2 embeddings (with a rejection threshold of 0.6).
Baseline Comparisons. We consider the following pre-training settings:
  • LmLm (Ours): Pre-trained on our annotated data with lookup calls, using the loss in Equation (1).
  • Standard: Pre-trained on our data without lookup calls. All other settings are identical.
  • Off-the-shelf Models: Models with publicly available pre-training weights, including: OpenAI/GPT2-124M, OpenAI/GPT2-355M, OpenAI/GPT2-774M [46], Pythia-1B [3], LLaMA2-7B [14], and LLaMA3.1-8B [13]. OpenAI/GPT2-124M and OpenAI/GPT2-355M are comparable in size to our LmLm models. These models are marked with an asterisk (*) in the results tables.
3.2 Learning to Lookup Facts is Easier than Memorization
Evaluation Setup. We first evaluate our models using language modeling perplexity on a held-out Wikipedia validation set. As LmLm introduces lookup calls, its perplexity is not directly comparable to that of Standard. To ensure a fair comparison, we report three variants of perplexity:
  • Static (Oracle): Assumes perfect lookup behavior, where LmLm always generates correct lookup calls and retrieves the correct values. This provides an optimistic lower bound. Perplexity is computed over all tokens excluding the lookup calls.
  • Dynamic: Reflects actual model behavior during inference. Lookup calls are generated and executed in real time, capturing failures due to incorrect queries or failed retrievals. Perplexity is again calculated over all tokens except the lookup calls.
  • Normalized: Measures the combined likelihood of generating the correct queries and subsequent text. Perplexity is computed over all tokens except for the retrieved values, and normalized by the number of tokens in the original unannotated text. See Appendix A.3 for formal definitions.
Figure 5: Validation perplexity comparison between Standard and LmLm on three variants of perplexity. Lower perplexity indicates better performance.
Perplexity Results. Figure 5 reports validation perplexities for both LmLm and Standard models.
If LmLm was benefiting from retrievals without learning to query correctly, we would expect Static perplexity to drop while Dynamic and Normalized worsen. Improvements across all variants indicate that LmLm is learning both to query and to generate more effectively.
We observe that LmLm consistently achieves lower perplexity than Standard across all model sizes and perplexity variants. In particular, LmLm achieves an average perplexity reduction of $1.98$ points under the Dynamic setting, demonstrating its effectiveness even with imperfect lookup calls. The Normalized variant highlights that LmLm assigns higher likelihoods to both lookup queries and grounded text, indicating improved training efficiency. These results support a key insight: learning to lookup specific facts is easier than memorizing them. Lastly, the observed gap between Dynamic and Static perplexities highlights potential for improvements to annotation quality and retrieval method.
Table 1: Evaluations on factual precision. FactScore evaluates open-ended biography generation with factual precision, while T-REx tests factual statement completion. Subscripts indicate the absolute difference compared to their respective Standard baselines.
FactScore (Open-ended Generation) T-REx (Knowledge Completion)
Model Model Type FActScore ↑ FActScore w/o len. penalty ↑ # Facts / Response ↑ Exact Match ↑ Precision @1 ↑
OpenAI/GPT2-124M - 14.6 14.7 24.2 20.1 20.3
GPT2-124M Standard 10.7 11.3 36.7 41.2 35.8
LmLm 20.6+9.9 20.9+9.6 43.4 54.6+13.4 40.3+4.5
LLaMA2-176M Standard 10.1 11.8 23.7 46.3 30.6
LmLm 30.6+20.5 32.4+20.6 31.4 54.1+7.8 34.5+3.9
OpenAI/GPT2-355M - 15.2 15.2 24.6 28.4 29.6
GPT2-355M Standard 14.4 14.4 38.9 44.9 38.9
LmLm 23.9+9.5 25.0+10.6 40.4 58.7+13.8 44.0+5.1
LLaMA2-382M Standard 14.0 16.6 32.4 52.0 34.5
LmLm 31.9+17.9 32.7+16.1 33.7 58.1+6.1 34.6+0.1
OpenAI/GPT2-774M - 17.4 17.4 49.4 35.6 31.9
Pythia-1B - 21.1 21.1 50.6 47.8 37.9
LLaMA2-7B - 34.0 35.1 33.9 60.5 44.4
LLaMA3.1-8B - 40.3 40.5 38.5 67.3 60.6
* Models marked with an asterisk (*) are off-the-shelf models with no additional training.
Downstream NLU performance. We compare LmLm and Standard models on five standard natural language understanding (NLU) tasks (Table 9). This evaluation serves as a sanity check to ensure that separating factual knowledge during pretraining does not come at the expense of general language understanding. We find that LmLm performs on par with Standard across all tasks, confirming that factual offloading preserves the model’s general-purpose capabilities.
3.3 Externalizing Knowledge Improves Factual Precision
FactScore Results. We evaluate factual precision on an open-ended biography generation task using prompts such as “Tell me a bio of Kang Ji-hwan. Kang Ji-hwan is”, and assess model outputs using the FactScore metric [36]. FactScore extracts atomic facts from generated text and measures the proportion that are supported by a trusted knowledge source.
Table 1 shows that LmLm substantially improves factual precision over Standard across all models. In particular, it outperforms Standard by $+20.5\%$ on LLaMA2-176M and $+17.9\%$ on LLaMA2-382M, demonstrating that even small models can produce factually accurate outputs when equipped with lookup capabilities. Remarkably, its performance approaches that of much larger models like Pythia-1B and LLaMA2-7B, despite having far fewer parameters.
T-REx Results. Additionally, we evaluate using the T-REx subset of the LAMA benchmark [45], which contains 11,615 samples. The task is to complete a short statement with a missing fact (e.g., a date or a place). For example, the statement “Jaideep Sahni (born 1968) is an Indian [MASK]” should be completed with the word “actor”. Following Schick et al. [50], we report Exact Match, which checks if the correct word appears within the first five predicted words, and Precision @1, which measures whether the model’s first predicted token matches the first token of the correct answer.
In Table 1, we observe that LmLm achieves large improvements on T-REx across all models. Specifically, it achieves an average gain of $10.3\%$ on Exact Match, demonstrating LmLm’s advantage on knowledge-intensive tasks. As our models are trained solely on Wikipedia, direct comparison to off-the-shelf models trained on broader web corpora is not entirely equitable. Nonetheless, the comparison between LmLm and Standard remains valid, as both share the same training data.
Table 2: Comparison of RAG vs. LmLm on factual precision. Results are shown for GPT2-355M.
Model FactScore (%) ↑ T-REx EM (%) ↑
Standard 15.2 28.4
Standard + RAG 20.1-3.8 75.8+17.1
LmLm 23.9 58.7
How Does LmLm Compare to RAG? While both LmLm and retrieval-augmented generation (RAG) incorporate external knowledge, their motivations differ fundamentally. RAG retrieves documents at inference time and prepends them to the model input [29, 23], requiring the LM to extract relevant information implicitly. In contrast, LmLm is trained end-to-end to issue structured lookup calls only when factual retrieval is needed, interleaving retrieval and generation at the entity level. This makes LmLm more precise and efficient: it avoids unnecessary retrievals, reduces token overhead, and discourages memorization of externalized facts. By contrast, RAG often retrieves irrelevant content and may inject information even when none is needed. That said, RAG is better suited to settings where relevant knowledge is long-form or unstructured—such as open-domain QA, or multi-hop reasoning. These settings fall outside LmLm’s current scope. Finally, we note that RAG and LmLm are complementary—RAG can be applied on top of LmLm for further gains.
To understand how LmLm compares empirically, we implement a standard RAG baseline that retrieves the top-4 most relevant Wikipedia articles and prepends them to a standard LLM during generation. While RAG is not the focus of our study, we include it as a point of reference. As shown in Table 2, LmLm achieves higher scores on FactScore at the current model scale, indicating the effectiveness of offloading factual knowledge at the entity level. However, RAG performs better on T-REx, possibly because retrieving full Wikipedia passages—rather than entity—level facts-provides broader coverage for knowledge completion tasks. We observe that small RAG models often struggle to make effective use of retrieved context and can be distracted by irrelevant information. Nonetheless, RAG exhibits encouraging levels of factuality, especially with instruction-tuned models.
3.4 Machine Unlearning: LmLm Supports Instant Forgetting
Figure 6: Evaluation of Machine Unlearning. We compare LmLm with NPO on the TOFU benchmark. Unlike prior methods, LmLm performs unlearning without any additional training. (a–b) Forget quality vs. utility trade-off. LmLm achieves ideal forgetting ($p$-value $>0.05$) without sacrificing general utility. (c–d) LmLm retains knowledge outside the forget set, unlike other methods that degrade retain-set performance due to parameter entanglement. Y-axis labels use -- or $\uparrow$ to indicate retention or improvement. For metrics like Forget ROUGE, where lower isn’t always better, we mark the retain model’s performance as reference.
One natural benefit of decoupling knowledge from model parameters is that editing and unlearning are achievable through simple operations on the external memory, without compromising the model’s general capabilities. To verify this, we extend our evaluation to a standard machine unlearning benchmark, TOFU [33].
Evaluation Setup. TOFU evaluates unlearning efficacy in a privacy-sensitive setting, where the goal is to selectively forget a targeted subset of information (the Forget Set) while preserving performance on the Retain Set and maintaining general model capabilities. The objective is to produce an unlearned model that is statistically indistinguishable from a model trained solely on the Retain Set (referred to as the retain model). The benchmark consists of 200 synthetic author profiles, each containing 20 QA pairs. It evaluates two key aspects:
  • Model utility: The average of three metrics—ROUGE (answer quality), Answer Probability, and Truth Ratio (likelihood assigned to the correct answer over distractors)—measured on the Retain Set, Real Author Set, and World Facts Set.
  • Forget quality: The $p$-value of a statistical test comparing the unlearned model with the corresponding retain model to assess whether the targeted knowledge has been effectively removed.
We use LLaMa-3.2-1B-Instruct as the base model and compare against NPO [60], a state-of-the-art unlearning method. For LmLm, we perform unlearning by simply removing entries in the database corresponding to the Forget Set. See Appendix B.2 for implementation details.
Unlearning Results. Figure 6 presents the results on the TOFU benchmark, where the Forget Set is 5% of the data. In Figure 6 (a-b), we show the forget quality and model utility throughout the unlearning process, where ideal performance is defined by effective forgetting without degrading model utility. LmLm achieves precisely this—effective forgetting with no loss in model utility—a direct benefit of decoupling factual knowledge from model parameters. Figure 6 (c–d) shows that LmLm preserves knowledge outside the Forget Set, whereas previous training-based methods tend to forget related information due to parameter entanglement. Importantly, forgetting is performed through simple database operations, without model updates or access to the Retain Set. In contrast, the strongest comparable baseline, NPO, incurs a utility degradation, while other unlearning methods either fail to forget or exhibit catastrophic drops in model utility (See Figure 4, right). While LmLm incurs upfront costs for data annotation and model training, it provides substantial payoff in use cases where knowledge editing, removal, or compliance with data deletion requests are necessary.
4 Discussion
Toward Efficient Scaling via Knowledge Offloading.
We summarize our main findings in Figure 4: (1) during pretraining, LmLm consistently obtains lower perplexity compared to its Standard counterparts. (2) LmLm achieves higher factual precision without compromising NLU performance. (3) LmLm supports instant forgetting while maintaining general model capabilities. In particular, even smaller instances of LmLm match or exceed the performance of much larger off-the-shelf models. These results highlight the potential of LmLm to scale efficiently by offloading knowledge to an external database, thereby maintaining strong factual accuracy with fewer parameters. While our experiments are limited to modest scales due to computational constraints, the observed trends suggest that LmLms offer a promising direction toward parameter-efficient language models that externalize factual storage. Such models may enable real-time, verifiable knowledge updates and open up new possibilities for deploying LmLm in resource-constrained or fact-sensitive environments.
Does LmLm Still Memorize Facts in Its Parameters?
Figure 7: Training loss on return value tokens.
We provide preliminary evidence that LmLm reduces factual memorization through its training design. Using the TOFU synthetic trainset, we compare training objectives by tracking the loss on return value tokens—the factual answers intended to be retrieved rather than memorized.
As shown in Figure 7, models trained with a standard SFT objective quickly reduce the loss on these tokens, suggesting memorization in the model’s parameters. In contrast, LmLm, trained with the masked loss maintains a high loss throughout training, indicating that these facts are not stored internally. LmLm’s successful application in the machine unlearning benchmark (Sec. 3.4) further supports this finding. Additionally, we observe a notable performance gap in Table 3, where factual precision drops substantially when the external database is disabled—forcing the model to rely on its internal parameters. These findings suggest that editing the database offers a direct way to control what the model knows and forgets.
Beyond Entity-Level Knowledge Offloading.
Table 3: Disabling retrieval significantly reduces performance on both FactScore and T-REx; see Table 8 for the full comparison.
Model Type FactScore (%) ↑ T-REx EM (%) ↑
Standard 14.0 52.0
LmLm (w/o database) 12.8-19.1 38.5-19.6
LmLm 31.9 58.1
By leveraging knowledge triplets, LmLm focuses on separating entity-level factual knowledge from linguistic competency. Current research on probing [44, 21], evaluating [36, 55], and editing [35] internal knowledge in language models similarly concentrates on entity-level atomic facts. However, extending our method beyond entity-level knowledge remains a challenge. Under the LmLm framework, this difficulty arises in determining effective formats for separating knowledge beyond entity-focused triplets. For instance, using a simple QA function call format raises concerns about the potential for hallucinated facts during annotation.
Additionally, while LmLm attempts to minimize the memorization of factual knowledge, some knowledge still remains unremoved. More comprehensive benchmarks for probing internal knowledge in language models, as well as distinct benchmarks for disentangling knowledge and reasoning, remain underexplored.
Limitations.
LmLm is a promising step toward separating factual knowledge from language models with many exciting future directions. Our current limitations include: (1) It does not guarantee perfect factuality during generation. Noise in the database and errors from fuzzy matching can introduce inaccuracies. However, such issues are easily traceable and verifiable for LmLm. (2) LmLm introduces additional tokens for lookup queries, which increases training and inference costs. (3) The current implementation focuses on entity-level factual knowledge, which captures only a subset of the broader factual spectrum. A more systematic framework is needed to determine which types of knowledge should be delegated to external databases. (4) Our experiments are limited to small models and datasets due to compute constraints. While sufficient to show core benefits, scaling up may improve performance and support more complex reasoning [17], which remains an open challenge for LmLm.
5 Related Work
Parametric vs Non-Parametric Knowledge.
Language models encode knowledge in two main forms: parametric, learned in model weights during training, and non-parametric, retrieved from external sources [34]. Early work showed that pre-trained models [11, 44] implicitly store factual knowledge. Scaling model size further improves factual recall [6, 24], but recent studies highlight fundamental limitations of parametric storage, including hallucination, staleness, poor attribution, and limited adaptability [42, 22, 26, 2], motivating the shift toward non-parametric approaches.
Retrieval-Augmented LLMs.
Non-parametric methods enhance language models with external sources to improve factuality and controllability. Retrieval-augmented generation (RAG)[29, 23, 52] retrieves relevant passages at inference time, while tool-augmented methods[50, 56] enable symbolic reasoning via external API calls. Retrieval-based pretraining approaches such as REALM[19] and RETRO[5] show that incorporating retrieval during training improves generalization and reduces memorization. More recently, SILO [37] demonstrates that even when excluding certain sources from pretraining, competitive perplexity can be maintained through proper retrieval mechanisms. Semi-parametric language models aim to decouple knowledge storage from generation by attaching non-parametric memories of key-value pairs. A prominent example is kNN-LM [27], which retrieves nearest-neighbor training examples at test time to adjust token probabilities without retraining. Follow-up works extend this idea to various settings [43, 28, 51, 57, 40]. Our approach builds on this line of work by focusing on scalable factual memory and dynamic usage during pretraining. Rather than storing full documents or relying on retrieval pipelines at inference time, our method enables the model to learn when to issue structured lookup calls and how to integrate the results into generation.
Machine Unlearning.
Machine unlearning aims to remove specific knowledge from a trained language model without full retraining [25, 7]. The goal is to forget a targeted subset of information while preserving the model’s general capabilities. Existing approaches fall into several main families. Gradient Ascent (GA) increases the loss on the forget set, while Gradient Difference (GD) jointly preserves performance on a retain set [32]. Preference-based methods, such as IdkDPO [33], NPO [60], and SimNPO [15], encourage the model to respond with neutral or deflective answers [47]. Among these, NPO achieves the strongest results on the TOFU benchmark [33], which evaluates entity-level forgetting and utility retention on a synthetic dataset of author profiles.
Despite strong results on TOFU, existing methods struggle to scale to larger forget sets (e.g., 5% of training data) and often degrade model utility or erase unintended related knowledge. These limitations arise from the entanglement of factual content and general linguistic ability in model parameters. Our approach takes a different route by structurally decoupling factual knowledge from the model. By offloading facts to an external database, forgetting reduces to removing entries—offering a scalable, verifiable, and precise solution without compromising utility. Moreover, LmLm can also be applied to knowledge editing benchmarks [38, 53, 16], where modifying or deleting facts becomes as simple as updating the database.
6 Conclusion
We introduce LmLms, a new class of language models for externalizing knowledge alongside an integrated solution to achieve this. Our results demonstrate promising trends towards efficient use of model capacity and offloading facts onto an external database. LmLm represents an alternative way to store facts during pre-training, and has the potential to be integrated with other common approaches developed for LLMs, including retrieval-based methods, symbolic reasoning, as well as knowledge representation. Consequently, LmLm opens up new ways for future language models to leverage the benefits of external knowledge databases such as verifiable updates – fundamentally, it is much easier and more memory-efficient to learn how to look up facts rather than to remember them.
Acknowledgments
SZ is supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. D24AP00259-00. CKB is supported by the National Science Foundation (NSF) through the NSF Research Traineeship (NRT) program under Grant No. 2345579. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). This work is partially supported by LinkedIn-Cornell Partnership, Open Philantropy, the National Science Foundation NSF (IIS-2107161, and IIS-1724282, HDR-2118310), the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875), DARPA, arXiv, and the New York Presbyterian Hospital.
Appendix A Implementation Details
A.1 Knowledge Extraction Details
Table 4: Prompt templates used for GPT-4o and Annotator annotation.
Model Prompt Template
GPT-4o You are a knowledge base construction expert. Extract entity-based factual knowledge from a passage and annotate it using the format: [dblookup(’Entity’, ’Relationship’) -> Value]. These annotations simulate a knowledge base query for factual generation. Place dblookup right after the entity and relationship appear, keeping the text flow natural.
Entity-Based Factual Knowledge Principles:
 - Entities: Use full names for people, organizations, places, or works.
 - Relationships: Use specific, reusable labels that define the connection clearly.
 - Values: Keep them concise and factual.
Annotation Principles:
 1. Extract ALL Atomic Facts: Each annotation should capture a single verifiable fact.
 2. Precise Annotations: Use correct and specific entity-relationship-value triples.
 3. Ensure Reusability: Use standardized and reusable entity and relation names.
 4. Contextual Positioning Rule: Place annotations only after both entity and relation appear.
 5. Preserve Text and Maintain Flow: Do not alter or disrupt the original text.
Example Annotation:
Input: Beyoncé Giselle Knowles-Carter (born September 4, 1981) is an American singer, songwriter, record producer, and actress.
Output: Beyoncé Giselle Knowles-Carter (born [dblookup(’Beyoncé Giselle Knowles-Carter’, ’Birth Date’) -> September 4, 1981] September 4, 1981) is an [dblookup(’Beyoncé Giselle Knowles-Carter’, ’Nationality’) -> American] American [dblookup(’Beyoncé Giselle Knowles-Carter’, ’Occupation’) -> singer, songwriter, record producer, actress] singer, songwriter, record producer, and actress.
Annotator Your task is to extract and annotate entity-based factual knowledge from the provided text. Identify and annotate specific entities, relationships, and values using the dblookup format: [dblookup(’Entity’, ’Relationship’) -> Value]
Annotation Guidelines:
 - Inline Insertion: Insert dblookup before factual statements without altering the text.
 - Atomic Facts: Each dblookup should capture one verifiable fact.
 - Entities: Use full names for people, organizations, places, or works.
 - Relationships: Use specific, reusable labels (avoid vague terms).
 - Values: Keep them concise and factual.
Model and Data.
We construct a high-quality seed dataset by sampling 1,000 passages from SQuAD-v2 [48] and 1,000 passages from Wikipedia. These passages are annotated by GPT-4o with structured factual triples and used to train the Annotator model. The remaining Wikipedia passages are used as the pre-training corpus for LmLm, with no overlap with the validation set.
Both the Corrector and Annotator are based on LLaMA-3.1-8B-Instruct, selected for its strong instruction-following capabilities. The Corrector and Annotator use a maximum context length of 2048 tokens. All input sequences are truncated to 1024 tokens during LmLm pretraining for consistency.
GPT-4o and Annotator annotations follow the format [dblookup(’Entity’, ’Relation’) -> Value]. These are converted to a token-based format for pretraining: <|db_start|> Entity <|sep|> Relation <|db_value|> Value <|db_end|>.
  • The Corrector is fine-tuned using LoRA (r=32, alpha=16[20] for 2 epochs on 19k GPT-4o-annotated SQuAD-v2 passages, with a learning rate of $2\times 10^{-4}$ and an effective batch size of 32. Sequence packing is enabled to improve training efficiency.
  • The Annotator is tuned with instruction on the 2k annotated passages using LoRA (r=32, alpha=16), with a learning rate of $2\times 10^{-4}$ and an effective batch size of 32. Training runs for 10 epochs with a maximum sequence length of 2048.
Annotation Pipeline.
We adopt a three-stage pipeline to distill GPT-4o’s structured annotations into a lightweight and scalable Annotator model:
  • Stage 1: Seed Annotation. GPT-4o is prompted to annotate input passages with structured factual triples (entity, relation, value). These are embedded directly into the text using the lookup call format. Prompt template is provided in Table 4.
  • Stage 2: Cleaning. A warm-start Corrector model is trained on the annotated data for 2 epochs without instruction prompts. Although underfit, it is effective at identifying noisy or ill-formed annotations. Specifically, we discard lookup calls where the token-level loss on the entity or relation is in the top 10% of the distribution.
  • Stage 3: Annotation. We instruction-tune a new Annotator model on the cleaned dataset for 10 epochs. This model learns to detect when factual knowledge should be externalized and how to issue structured lookup queries. The trained model is then applied to the full pre-training corpus to generate large-scale factual supervision. Prompt template is provided in Table 4.
Additional Notes.
We observe a bimodal loss distribution on entity and relation tokens in the GPT-4o-generated annotations. This is likely due to GPT-4o accessing future context during generation, which breaks the left-to-right constraint of autoregressive models. As a result, some annotations are not recoverable from preceding context alone.
The Corrector helps filter out such cases—removing lookup calls that are (1) not inferable from prior context, (2) overly specific or inconsistent, or (3) syntactically malformed. This filtering improves the quality of supervision provided to the final Annotator model.
Ultimately, the Annotator learns to insert lookup calls only when they are contextually grounded and likely to enhance factual accuracy. This encourages retrieval-based reasoning and helps LmLm offload factual knowledge from its parameters into a structured database.
A.2 Model Architecture and Training Details
We pretrain LmLm from scratch using GPT-2 and LLaMA2-style decoder-only architectures. Each model uses its original tokenizer and vocabulary, extended with four special tokens reserved for lookup calls. This results in a vocabulary size of 50,261 for GPT-2 models and 32,004 for LLaMA2 variants. Full architecture specifications, including hidden size, depth, and parameter counts, are shown in Table 5.
All models are trained for 8 epochs with a context length of 1,024 tokens using mixed-precision training. For LLaMA2-176M and LLaMA2-382M, we use a batch size of 256 and train for 105k steps, totaling approximately 8 H100-days. Training is performed using Hugging Face Accelerate in bf16 precision. Hyperparameters such as learning rate, scheduler, and warmup steps are detailed in Table 6.
Table 5: Model architecture, vocabulary size (including 4 special tokens), and parameter counts. We report both total and non-embedding parameter counts.
Model Hidden Size #Layers #Heads Vocab Size Params (Total / Non-Embed)
GPT2-124M 768 12 12 50,261 124.4M / 85.5M
LLaMA2-176M 512 8 8 32,004 176.4M / 160.0M
GPT2-355M 1024 24 16 50,261 354.8M / 303.4M
LLaMA2-382M 768 12 12 32,004 381.8M / 357.3M
Table 6: Training hyperparameters. LLaMA2-176M, LLaMA2-382Mare initialized in $float32$ and trained with mixed precision ($bf16$) using Hugging Face Accelerate.
Model Batch Size Total Steps LR Scheduler Warmup Precision
GPT2-124M,GPT2-355M 320 66k 5.0e-4 $bf16$ (mixed)
LLaMA2-176M, LLaMA2-382M 256 105k 5.0e-4 cosine 2000 $bf16$ (mixed)
A.3 Formalization of Training and Evaluation Objectives
We denote an autoregressive language model by $p_{\theta}$, which defines a probability distribution over a sequence of tokens $x=(x_{1},\dots,x_{T})$ as:
$p_{\theta}(x)=\prod_{t=1}^{T}p_{\theta}(x_{t}\mid x_{<t}).$
Each token $x_{t}$ belongs to one of the following categories:
  • $\mathcal{T}_{\text{org}}$: original text tokens from the raw corpus;
  • $\mathcal{I}_{\text{train}}$: Return values and <|db_end|> are excluded from training loss objective.
  • $\mathcal{T}_{\text{e}}$, $\mathcal{T}_{\text{r}}$: tokens representing entities and relation arguments within database lookup calls;
  • $\mathcal{T}_{\text{v}}$: tokens corresponding to retrieved factual values (i.e., return values);
  • $\mathcal{T}_{\text{db}}$: special tokens used to mark database lookup segments, including:
    • <|db_start|>: begins a lookup call;
    • <|sep|>: separates entity and relation in the query;
    • <|db_retrieve|>: signals the insertion point for the returned value;
    • <|db_end|>: marks the end of the lookup block.
Here is an example using background color to highlight different token categories:
$\mathcal{T}_{\text{org}}$:
Napoleon was born on <|db_start|> Napoleon <|sep|> Birth_Date <|db_retrieve|> August 15, 1769 <|db_end|> August 15, 1769.
$\mathcal{I}_{\text{train}}$:
Napoleon was born on <|db_start|> Napoleon <|sep|> Birth_Date <|db_retrieve|> August 15, 1769 <|db_end|> August 15, 1769.
Training Loss.
The training objective excludes supervision over return values and the closing marker <|db_end|> to prevent memorization of factual knowledge:
[…]
Evaluation Metrics.
We report both perplexity and negative log-likelihood (NLL), computed over different token subsets depending on the evaluation setting:
  • Static & Dynamic Perplexity: Tokens corresponding to lookup calls and return values are excluded:
    $\text{PPL}_{\text{static/dynamic}}=\exp\left(-\frac{1}{|\mathcal{I}_{\text{org}}|}\sum_{t\in\mathcal{I}_{\text{org}}}\log p_{\theta}(x_{t}\mid x_{<t})\right).$
  • Normalized Perplexity: This metric fairly compares generation likelihood by excluding retrieved factual values from the loss, but normalizing by the length of the original (fully reconstructed) text. Specifically:
    $\text{PPL}_{\text{norm}}=\exp\left(-\frac{1}{|\mathcal{T}_{\text{org}}|}\sum_{t\in\mathcal{I}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t})\right)$
  • Negative Log-Likelihood (NLL): Matches the training loss computation:
    $\text{NLL}(x)=-\sum_{t\in\mathcal{I}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t}).$
A.4 Database and Retrieval Setting
We build our database by annotating the pre-training data, obtaining 54.6M knowledge triplets consisting of 9.5M unique entities, 8.5M relationships and 16.2M unique values. For retrieval, we employ a fuzzy matching mechanism based on the cosine similarity of sentence embeddings from all-MiniLM-L6-v2333https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. As shown in Table 8, specifically, given a lookup call, we compute its embedding and compare it with the embeddings of stored triplets in our database. If the highest similarity score is below a threshold of 0.6, we return unknown to indicate that no sufficiently similar match was found. Alternatively, we implement a prefix-tree constrained generation, ensuring that lookup calls remain covered by the structured knowledge representations. See detailed discussion in Appendix D.3.
Figure 8: Unconstrained vs. Constrained Query Generation.
Appendix B Experimental Setup
B.1 Evaluation Benchmarks
Perplexity.
We evaluate language modeling perplexity on a held-out Wikipedia test set consisting of 1,000 passages (~245k tokens). We use the same tokenizer as the original model (either GPT2 or LLaMA2) and apply Hugging Face’s Trainer with sequence packing enabled. For LmLm, reference completions are annotated using Annotator. In the dynamic setting, the model generates its own lookup arguments, but we force lookup calls to occur at the same positions as in the reference. The exact formulation of perplexity is provided in Appendix A.3.
FactScore.
We evaluate factual precision using FactScore [36], a benchmark for open-ended biography generation. Given a generated text, FactScore extracts a set of atomic facts and computes the proportion that is supported by a trusted knowledge source. We use the first 100 biography queries provided in the benchmark. All models generate outputs using greedy decoding (maximum length = 256 tokens; repetition penalty = 1.2). Factuality is validated using retrieval-augmented prompting with ChatGPT, following the official evaluation protocol. [^4] https://github.com/shmsw25/FActScore
For LmLm, which is not instruction-tuned, we use a fixed prompt template to elicit biography completions: “Tell me a bio of <name>. <name> is” This prompt is applied consistently across all samples. To encourage structured queries during generation, we apply a logit bias to four special tokens in the vocabulary: <|db_start|>, <|sep|>, <|db_retrieve|>, and <|db_end|>, with respective bias values of 5.0, 2.0, 2.0, and 2.0. Retrieval is performed using fuzzy matching with a cosine similarity threshold of 0.6. If no relevant triplet is found, the model continues generation using the plain text string unknown as a fallback. This behavior is untrained and left for future work to improve robustness to retrieval failures.
T-REx.
We adapt the T-REx subset from LAMA [44] for autoregressive models by filtering examples in which the masked entity does not appear in the final position. This yields 11,615 left-to-right compatible examples, following Schick et al. [50]. Each input consists of a factual statement, such as: “Jaideep Sahni (born 1968) is an Indian [MASK]” The model is expected to complete the statement with a single token (e.g., actor). We evaluate using two metrics: Exact Match, which checks whether the reference answer appears among the first five generated content words, and Precision@1, which checks whether the first generated content token matches the reference.
All models generate outputs using greedy decoding, and both metrics are computed after post-processing to remove lookup calls. For LmLm, we enforce a database lookup call at the masked position using fuzzy matching with a similarity threshold of 0.6. If no match is found, the model continues with standard decoding without triggering a structured lookup call, as the target fact may belong to common knowledge not covered by the database.
B.2 Machine Unlearning Setting
TOFU.
The TOFU benchmark [33] evaluates unlearning efficacy in privacy-sensitive scenarios, aiming to selectively remove a specific subset of information (the Forget Set) from a model (the Full Model) while maintaining performance on retained information (the Retain Set) and general model capabilities. The benchmark’s primary goal is to ensure the resulting Unlearned Model is statistically indistinguishable from a model trained exclusively on the Retain Set (the Retain Model).
The TOFU benchmark comprises 200 synthetic author profiles, each associated with 20 QA pairs. Performance is assessed using two primary metrics:
  • Model utility: An average of three metrics—ROUGE (answer quality), Answer Probability (the likelihood of correct answers), and Truth Ratio (the likelihood assigned to correct answers over distractors)—evaluated on the Retain Set, Real Author Set, and World Facts Set.
  • Forget quality: The $p$-value from a statistical test comparing the unlearned model to the corresponding retain model, quantifying the effectiveness of knowledge removal.
Our implementation directly builds upon the official TOFU repository [^5] https://github.com/locuslab/open-unlearning.
We use LLaMa-3.2-1B-Instruct as the base model and compare against NPO [60], a state-of-the-art method for unlearning. We test on forget 5% setting. For LmLm, unlearning is implemented simply by removing relevant entries corresponding to the Forget Set from the external database.
NPO baseline.
Negative Preference Optimization (NPO) [60] is a state-of-the-art method for selective knowledge removal, especially effective for unlearning large portions (50%-90%) of training data. Unlike traditional gradient-based unlearning methods, which often degrade a model’s general performance and struggle to forget only 10% of training data, NPO explicitly guides the model away from undesired (negative) samples, maintaining stable training dynamics and preventing catastrophic performance degradation. Further details can be found in the original paper [60]. We follow the official TOFU implementation, running unlearning fine-tuning three times with different random seeds and reporting the mean and variance in Figure 6.
LmLm Implementation of TOFU Evaluation.
To evaluate unlearning effectiveness, we assume the synthetic knowledge used in TOFU is fully represented in our external database. We annotate the complete TOFU dataset (4k synthetic QA pairs) using GPT-4o, subsequently building our database from these annotations.
Since LmLm is applied to pre-existing models, we introduce an additional step to ensure the model can utilize the lookup mechanism effectively. Specifically, we first perform a warm-up training stage on annotated Wikipedia data, followed by fine-tuning on the annotated TOFU training set, using the same hyperparameters as the baseline models (see Table 7).
For ROUGE evaluations, generated answers are post-processed to remove structured lookup tokens before computing the scores. For likelihood-based metrics (Answer Probability and Truth Ratio), we evaluate the model’s probabilities only on the annotated training input segments (_$\mathcal{I}\_\text{train}$, as defined in Appendix A.3). Thus, ROUGE scores remain directly comparable across methods, but likelihood-based metrics are not directly comparable due to differences between raw and annotated reference answers.
Interpreting TOFU Results.
When interpreting TOFU results, the key indicator of successful unlearning is how closely the unlearned model’s performance matches that of the retain model. Ideally, the two models should be indistinguishable, indicated by a forget quality $p$-value above 0.05. In practice, this means the unlearned model should achieve forget quality above 0.05 while maintaining consistent model utility, and its ROUGE scores on both the Forget Set and Retain Set should closely resemble those of the retain model. Importantly, lower ROUGE scores on the Forget Set do not necessarily imply better unlearning, as these reductions could also result from general performance degradation.
B.3 Experimental Settings by Figure
We detail the experimental configurations corresponding to each figure in the main paper:
  • Figure 4: (Left) We eval on a held-out wikipedia validaition set of 100 passages (~21k tokens) every 1000 steps during pretraining. (Middle) The detailed results for FactScore and NLU are in Table 1 and  9. (Right) The backbone model is LLaMA-3.2-1B-Instruct. We implement LmLm evaluation based on TOFU official repo It is forget 5% setting. The detailed results for baselines methods are copied from TOFU repo.
  • Figure 6: For NPO baselines, we follow the official TOFU implementation, running unlearning fine-tuning five times with different random seeds (0, 42, 420, 69, 4497) and reporting the mean and variance throughout training. The ideal performance of the retain model is indicated with a marker. For LmLm, which does not require training to unlearn, we show only the pre- and post-unlearning results. Details on how to interpret the results are in Appendix B.2.
  • Figure 7: We use the TOFU synthetic training set to compare training objectives by tracking the loss on return value tokens—the factual spans intended to be retrieved via lookup. Models are evaluated every 125 steps during training.
Table 7: Training hyperparameters used in different experimental settings. (Full set: 4k QA pairs, Retain set: 3.8k QA pairs)
Setting learning_rate warmup_steps num_train_epochs batch_size dataset
Finetune (Standard) 1e-5 0.2 5 32 TOFU trainset
Warmup (LmLm) 5e-5 0.25 1 64 Annotated Wikipedia (9.8k chunk)
Finetune (LmLm) 5e-5 0.2 5 32 Annotated TOFU trainset
Appendix C Detailed Results
RAG.
[…]
No fine-tuning is performed; models are evaluated with retrieval at inference time only. In practice, we observe that the BM25 retriever consistently retrieves highly relevant supporting documents, ensuring a fair and competitive RAG baseline. We evaluate RAG with OpenAI/GPT2-355M, enabling a direct comparison against our LmLm model of similar scale.
Factual Degradation When Forcing Internal Recall.
Table 8 shows the full results when database access is disabled, forcing models to rely solely on internal parameters. Across all variants, we observe a consistent drop in factual precision, often below the Standard baseline. This degradation supports our main claim: LmLm does not memorize factual answers but retrieves them externally. These results highlight that what the model knows and forgets is determined by the database, enabling precise and direct control through simple edits.
Table 8: Impact of database access on factual precision. Disabling access leads to performance drops in both FactScore and T-REx, confirming that LmLm relies on retrieval from external database rather than memorization.
Metrics
Model Model Type Database FActScore (%) ↑ T-REx Exact Match (%) ↑
GPT2-124M Standard - 10.7 41.2
LmLm $\times$ 14.9-5.7 32.0-22.6
LmLm $\checkmark$ 20.6 54.6
LLaMA2-176M Standard - 10.1 46.3
LmLm $\times$ 11.3-19.3 34.9-19.2
LmLm $\checkmark$ 30.6 54.1
GPT2-355M Standard - 14.4 44.9
LmLm $\times$ 10.4-13.5 36.4-22.3
LmLm $\checkmark$ 23.9 58.7
LLaMA2-382M Standard - 14.0 52.0
LmLm $\times$ 12.8-19.1 38.5-19.6
LmLm $\checkmark$ 31.9 58.1
Additional Loss Curves.
Figure 9 shows the full training loss curves on different token types in the TOFU synthetic trainset. While the main paper focuses on return value tokens, we include loss on $\mathcal{T}_{\text{org}}$, entity, and relationship tokens here for completeness. We observe no significant difference between training objectives on these tokens.
Figure 9: Training loss on $\mathcal{T}_{\text{org}}$ tokens, entity tokens, relationship tokens, and return value tokens.
Appendix D Further Analysis
D.1 Does LmLm Affect Language Understanding?
Beyond the promising results, it is important to verify that our approach does not compromise the general capabilities of pretrained language models. To assess this, we follow Penedo et al. [41] and evaluate on a set of "high-signal" Natural Language Understanding (NLU) benchmarks (as shown in Table 9). Given our focus on smaller models, we exclude benchmarks where both GPT-2 and similarly sized models fail to rise above the noise floor [12]. This leaves us with the following benchmarks: Commonsense QA [54], HellaSwag [58], PIQA [4], SIQA [49], and ARC Easy [10]. Implementation details can be found in lighteval [^7] https://github.com/huggingface/lighteval.
It is important to mention that Standard and LmLm are pretrained solely on the Wikipedia dataset, which makes certain benchmarks not applicable for comparison with off-the-shelf models. However, the chosen NLU benchmarks still effectively address concerns about the potential negative impact on the models’ general language understanding ability introducing by removing factual knowledge.
Table 9: Evaluation of NLU benchmarks using normalized accuracy metrics, demonstrating that separating factual knowledge during pretraining does not compromise overall model performance.
Metrics
Model Model Type CSQA HellaSwag PIQA SIQA ARC Easy All
Random Chance - 20.0 25.0 50.0 33.3 25.0 30.7
OpenAI/GPT2-124M - 30.3 29.8 62.5 40.7 39.5 40.6
GPT2-124M Standard 26.5 26.4 55.3 39.2 34.2 36.3
LmLm 27.9 26.8 55.1 39.9 35.0 37.0
LLaMA2-176M Standard 26.6 27.0 55.4 40.4 33.9 36.7
LmLm 26.8 28.2 55.2 40.2 35.8 37.2
OpenAI/GPT2-355M - 32.6 37.1 66.4 41.2 43.6 44.2
GPT2-355M Standard 28.1 27.0 55.7 40.0 37.8 37.7
LmLm 27.1 27.7 56.8 40.1 36.9 37.7
LLaMA2-382M Standard 27.8 28.8 55.2 41.0 35.8 37.7
LmLm 26.9 29.1 56.1 40.8 35.9 37.8
* Models marked with an asterisk (*) are off-the-shelf models with no additional training.
D.2 Entity Frequency Analysis in the Knowledge Database
Figure 10: Frequency distribution of database entities matched to Wikipedia entries. The database spans a wide range of entity frequencies, including many long-tail cases.
We analyze the prevalence of knowledge triplets in our database by estimating how often each entity appears in Wikipedia. Using entity frequency statistics from Kandpal et al. [26], we apply fuzzy string matching (threshold = 70) to align database entities with Wikipedia entries. Figure 10 shows the distribution of matched entity frequencies. The results indicate that our database spans both common and long-tail knowledge, with a substantial portion of entities being less frequent in the overall training corpus.
D.3 Ablation: Unconstrained vs. Constrained Query Generation
We compare fuzzy matching with prefix-tree constrained decoding for generating (entity, relation) queries. In the unconstrained setting, the model freely generates queries, which are then matched against the database using cosine similarity over sentence embeddings. This approach provides flexibility but may result in syntactically invalid or ambiguous queries.
In contrast, prefix-tree decoding restricts generation to valid entries encoded in a trie structure, ensuring syntactic correctness and reducing hallucinations. It is also compatible with beam search and nucleus sampling, allowing the model to explore multiple valid paths in the induced knowledge graph. See Appendix A.4 for more details.
As shown in Table 10, fuzzy matching consistently outperforms prefix-tree decoding in our setting, and is therefore used in all reported experiments. While prefix-tree decoding offers stronger structural guarantees, we find its diversity can be overly constrained by a relatively small database. As both the database and model scale, we expect the benefits of structured decoding to become more pronounced.
Table 10: Comparison of Fuzzy Match and Prefix-Tree Decoding. Fuzzy matching, used by default in LmLm, offers higher flexibility; prefix-tree decoding is included for ablation.
Model Model Type Decoding FActScore (%) ↑ FActScore w/o len. penalty ↑ Facts / Response ↑
LLaMA2-176M Standard - 10.1 11.8 23.7
LmLm Prefix-tree 23.0-7.6 23.9-8.5 34.6+3.2
LmLm Fuzzy Match 30.6 32.4 31.4
LLaMA2-382M Standard - 14.0 16.6 32.4
LmLm Prefix-tree 23.5-8.4 31.3-1.4 28.1-5.6
LmLm Fuzzy Match 31.9 32.7 33.7
D.4 Qualitative Results
To complement our quantitative results, we present qualitative examples comparing outputs from LmLm, Standard, and off-the-shelf models. As shown in Table 11,  12 and 13, LmLm produces concise, factually grounded responses by leveraging external knowledge, whereas standard and off-the-shelf models often include hallucinated content. These examples illustrate how knowledge offloading enables LmLm to maintain factual precision, further supporting our central finding (Figure 4) that LmLms achieve more with less by scaling efficiently through externalized knowledge.
Table 11: Qualitative examples of generated outputs from models with Standard and LmLm settings.
Model Generated Output
LLaMA2-176M-LmLm Ko Itakura is the <|db_entity|> Ko Itakura<|db_relationship|> Position<|db_return|> center-back, defensive midfielder<|db_end|> center-back and defensive midfielder for <|db_entity|> Ko Itakura<|db_relationship|> Current Team<|db_return|> Japan national team<|db_end|> Japan’s senior squad in the <|db_entity|> Ko Itakura<|db_relationship|> League Participation<|db_return|> 2022 FIFA World Cup<|db_end|> 2022 FIFA World Cup. He was born on <|db_entity|> Ko Itakura<|db_relationship|> Birth Date<|db_return|> January 27, 1997<|db_end|> January 27, 1997. Career. On <|db_entity|> Ko Itakura<|db_relationship|> Joined Club<|db_return|> Kawasaki Frontale<|db_end|> July 30, 2015, it was announced that Itakura joined JFL club Kawasaki Frontale from their youth system. In <|db_entity|> Ko Itakura<|db_relationship|> Moved To<|db_return|> Groningen<|db_end|> August 2016 he moved to Groningen. On <|db_entity|> Ko Itakura<|db_relationship|> Debut Date<|db_return|> June 17, 2019<|db_end|> June 17, 2019, Itakura debuted against <|db_entity|> Ko Itakura<|db_relationship|> Debut Opponent<|db_return|> Uruguay<|db_end|> Uruguay in an international friendly match. After playing only one game for Groningen during his debut season, Itakura left the club at the end of the year.
LLaMA2-176M-Standard Ko Itakura is the name given to an ancient Japanese settlement in what is now part of Shimane Prefecture, Japan. The site was discovered by archaeologists from Kyoto University and has been excavated since 1978. The first written reference to this place appears in "Nihon Kōki", which dates back to the year 604 AD (Kumano no Kan’i), but it seems that there were already settlements at this location before then. In the early Heian period, the area around modern-day Shimonoseki City became known as , meaning "the land between rivers". During the Kamakura period, the region came under control of the Ashikaga clan who built their capital on Mount Hiei near present-day Naka Ward. After the fall of the Ashikaga shogunate, they lost control over much of eastern Japan, including the area surrounding modern-day Shimane prefecture. However, after the Battle of Sekigahara, Tokugawa Ieyasu established himself as lord of most of northern Japan with his domain centered on Matsue Castle. This castle later fell into disrepair during the Edo Period, and
Table 12: Qualitative examples of generated outputs from models with Standard and LmLm settings.
Model Generated Output
LLaMA2-382M-LmLm Ko Itakura is the <|db_entity|> Ko Itakura<|db_relationship|> Position<|db_return|> center-back, defensive midfielder<|db_end|> center-back and defensive midfielder for <|db_entity|> Ko Itakura<|db_relationship|> Team<|db_return|> Japan national team<|db_end|> Japan’s national team. He was born in <|db_entity|> Ko Itakura<|db_relationship|> Birthplace<|db_return|> Yokohama<|db_end|> Yokohama on <|db_entity|> Ko Itakura<|db_relationship|> Birth Date<|db_return|> January 27, 1997<|db_end|> 27th January 1997. His father is from <|db_entity|> Ko Itakura<|db_relationship|> Father’s Origin<|db_return|> Japanese<|db_end|> Japanese and his mother is from <|db_entity|> Ko Itakura<|db_relationship|> Mother’s Origin<|db_return|> Japanese<|db_end|> Japanese. When he was young, he played baseball but after watching an exhibition match against a professional baseball club, he decided to become a footballer. In <|db_entity|> Ko Itakura<|db_relationship|> Joined Club Year<|db_return|> Kawasaki Frontale<|db_end|> 2013, he joined J1 League side Kawasaki Frontale. However, he could not play many matches behind Shusaku Nishikawa until <|db_entity|> Ko Itakura<|db_relationship|> First Match as Starter<|db_return|> Uruguay<|db_end|> September when he debuted at right back against Uruguay. After that, he became a regular player under manager <|db_entity|> Ko Itakura<|db_relationship|> Manager Under Whom Became Regular Player<|db_return|> Japan national team<|db_end|> Shinji Ono. On <|db_entity|> Ko Itakura<|db_relationship|> Debut Date<|db_return|> June 17, 2019<|db_end|> 17 June 2019, he debuted with Japan national team against Chile during the <|db_entity|> Ko Itakura<|db_relationship|> Competition Debut<|db_return|> Uruguay<|db_end|> 2019 Copa América. Career statistics. “Updated to end of 2018 season”. National team career. In August 2016, Itakura was elected to the <|db_entity|> Ko Itakura<|db_relationship|> U-23 Selection<|db_return|> Japan U-20 national team<|db_end|> Japan U-20 national team for the <|db_entity|> Japan U-20 national team<|db_relationship|> Tournament Participation<|db_return|> 2017 U-20 World Cup<|db_end|> 2017 U-20 World Cup. At this tournament, he played all 4 matches as left back of three back defense. In May 2019, he was selected for the <|db_entity|> Ko Itakura<|db_relationship|> Senior Squad Selection<|db_return|>
LLaMA2-382M-Standard Ko Itakura is the first album by Japanese singer-songwriter Koda Kumi, released on November 19, 2003 under Sony Music Japan label. The album was reissued in March 2004 with two bonus tracks: “Kimi no Uta” and “Tell Me A Bio”.
Table 13: Qualitative examples of generated outputs from different off-the-shelf models.
Model Generated Output
OpenAI/GPT2-124M Ko Itakura is the author and editor of The Art Of Anime, which has been published by Kadokawa since 2013 (and was also translated into English in 2015). I’m not sure if I can say much more about this book than what you’ve said before but it’s pretty good! This manga really does have an interesting story arc that makes for some great reading experience as well as being very entertaining to read on its own terms!!
OpenAI/GPT2-355M Ko Itakura is an American author, illustrator and animator who has worked on many anime series including: Naruto, One Piece (one of the most popular manga in Japan), Bleach, Dragon Ball Z, Ghost in The Shell, Gintama.
He also created characters for various games such as Final Fantasy VII, Pokémon X/Y, or Pokemon Omega Ruby/Alpha Sapphire.
The story takes place during World War II when Japanese soldiers are sent to fight against Nazi Germany’s forces under General Yamamoto at Kure Castle near Tokyo.
After defeating them all except one—Genosha-Kaiji from Earth Kingdom fame—they return home with their families but find that there was no sign left behind by those killed fighting alongside them; instead it seems like everything had been destroyed!
Now this isn’t exactly what you’d expect if your favorite character were actually dead… but then again maybe he wasn’t?
Well I’m sure we’ll get into more details about him later so stay tuned…
Pythia-1B Ko Itakura is the name of my father.\̈newline Ï don’t know him,s̈aid the man, b̈ut I can tell you that he was an excellent swordsman and a good fighter. He had many friends in the army who were killed by bandits or robbers when they tried to capture them. They would have been glad if someone like him could help them escape from their captors. But it’s not possible for anyone with his skill to be captured alive. So we must find out where he lives so that we may rescue him.\̈newline The man went on: Ḧe has lived here all these years because there are no other people around. The only thing that keeps him going is the hope that someday some day he will return home. And now he wants to see you again. Please come this way.\̈newline They walked along the road until they came to a small house which looked as though it might belong to one of the villagers. A woman opened the door and invited them inside. She led them into a room full of bookshelves. There were several large tables covered with maps and charts. On each table sat a young girl about ten years old. Her hair was long and black and her eyes were bright blue. She wore a white dress and carried
LLaMA2-7B Ko Itakura is the author of *The Artist’s Way* (4.06 avg rating, 12 ratings, 3 reviews), *The Artist’s Way Workbook: A Course in Discovering and Recovering Your Creative Self* (5/5(7)).
*The Artist’s Way* by Julia Cameron – Book Review | The Guardian
Julia Cameron has been an active artist for more than thirty years. She was born in London but now lives in New York City with her husband, Martin. They have three children together. In addition to writing books on creativity, she also writes plays and musicals. Her latest book, *Finding Water*, will be published this fall.
She is currently working on another novel called *The Last Days Of Summer*. This article contains affiliate links which means that if you click through from here we may receive compensation at no extra cost to yourself! We only recommend products or services that we believe are useful for our readers so please feel free to check them out before making any purchases :) Thank you for reading our blog post today!! If there’s one thing I know about myself it’s how much I love reading blog posts like these ones because they always make my day better when someone
Table 14: Example of input and output from the RAG model. The model is prompted using retrieved documents and a user query.
Field Content
[…]
Retrieved Documents "Kang Ji-hwan" is a South Korean actor born Jo Tae-gyu on March 20, 1977. He began in musical theatre before transitioning to television and film, with notable roles in dramas like "Be Strong, Geum-soon!", "Capital Scandal", and "Hong Gil-dong".
"90 Days, Time to Love" is a 2006 South Korean TV series starring Kang Ji-hwan, centered on a professor and scriptwriter who rekindle a romance after discovering they are cousins.
Kang also produced and starred in the musical "Cafe In", becoming the first Korean actor to perform at Tokyo Globe Theatre.
User Query Tell me a bio of Kang Ji-hwan.
Model Input Answer the question or complete the prompt based on the given document. The following are given documents:
"Kang Ji-hwan" is a South Korean actor born Jo Tae-gyu on March 20, 1977. He began in musical theatre before transitioning to television and film, with notable roles in dramas like "Be Strong, Geum-soon!", "Capital Scandal", and "Hong Gil-dong".
"90 Days, Time to Love" is a 2006 South Korean TV series starring Kang Ji-hwan, centered on a professor and scriptwriter who rekindle a romance after discovering they are cousins.
Kang also produced and starred in the musical "Cafe In", becoming the first Korean actor to perform at Tokyo Globe Theatre.
Tell me a bio of Kang Ji-hwan.
Model Output (OpenAI/GPT2-355M) Kang Ji-hwan is known primarily for portraying characters with complex emotions like "the heartbreak," which can be seen throughout most drama shows nowadays. His character arc often involves both romantic relationships and difficult situations where people have conflicting feelings about their own lives. For example during this season we see how complicated it is for two men trying desperately not get divorced because there isn’t enough time left before graduation! What do you think? Do your favorite actors play these kinds/characters differently than others? Let us know what kind(ies?) I missed out below…
Model Output (Llama-3.2-1B-Instruct) Kang Ji-hwan is a renowned South Korean actor known for his versatility and range in various genres. Early Life Kang Ji-hwan was born on March 20, 1977, in Busan, South Korea. His full name is Jo Tae-gyu. Career Beginnings He started his career in musical theater before transitioning into film and television. Kang debuted in musical performances in 2001 with stints in productions like "The Rocky Horror Show" and "Grease". Breakthrough Roles His breakthrough came in 2005 with the drama "Be Strong, Geum-soon!". He later starred in "Capital Scandal", "Hong Gil-dong", and "90 Days, Time to Love", showcasing his ability to play complex characters. Film Career Kang ventured into cinema with movies like "Rough Cut" (2008) and "My Girlfriend Is an Agent" (2009). Recognition Throughout his career, Kang has received numerous awards and nominations, solidifying his status in Korean entertainment.